OpenStudy (anonymous):

what's the derivative of y=ln square root of 1+ sin2x

OpenStudy (anonymous):

\[y= \ln \sqrt{1+\sin2x}\]

OpenStudy (amistre64):

logs pull out exponents ... ln(a^b) = b ln(a) that should at least simplify the process

OpenStudy (amistre64):

take the "e" of both sides, that will undo the ln

OpenStudy (amistre64):

since the constant is trivial .... \[y= ln(x)\] \[e^y= e^{ln(x)}\] \[e^y= x\] \[y'e^y= x'\] \[y'= e^{-y}\]

OpenStudy (amistre64):

i feel like if made an error ....

OpenStudy (amistre64):

in this case, i was trying to signify x' as the derivative of the innards ... so x' not equal to 1. bad choice of variable name on my part :)

OpenStudy (anonymous):

why you took 'e'? no need to cancel out the ln .. can't you just do this y=ln (1+sin2x)^1/2

OpenStudy (amistre64):

yeah, i was thinking integration with that .... i blame these tired old eyes :/

OpenStudy (amistre64):

good catch tho ... \[y= \ln \sqrt{1+\sin2x}\] \[y= \frac12\ln (1+\sin2x)\]

OpenStudy (anonymous):

didn't get what you did..

OpenStudy (amistre64):

i just made it less scary by pulling out the sqrt of it.

OpenStudy (anonymous):

how it came 1/2 ln ...

OpenStudy (amistre64):

\[\sqrt{n}=n^{1/2}\] \[log(a^b)=b~log(a)\]

OpenStudy (amistre64):

these are algebraic propeties

OpenStudy (amistre64):

and the derivaitive of a log is:\[D[log(u)]=\frac{u'}{u}\]

OpenStudy (anonymous):

ohhhh this point:) thaanks:D

OpenStudy (anonymous):

you made it easy sir

OpenStudy (amistre64):

:) i do get there in the end

OpenStudy (anonymous):

hats off to you